OpenStudy (anonymous):

Determine two pairs of polar coordinates for the point (2, -2) with 0° ≤ θ < 360° (2 square root of 2, 225°), (-2 square root of 2, 45°) (2 square root of 2, 135°), (-2 square root of 2, 315°) (2 square root of 2, 315°), (-2 square root of 2, 135°) (2 square root of 2, 45°), (-2 square root of 2, 225°)

OpenStudy (anonymous):

@Hero @KingGeorge @Psymon ?

OpenStudy (psymon):

Yo, I have been summoned

OpenStudy (anonymous):

haha yes, again.

OpenStudy (psymon):

So first we need to translate this into polar coordinates. There are several conversions we need to be able to remember for these conversions back and forth \[x = rcos \theta \] \[y = rsin \theta \] \[x ^{2}+y ^{2} = r ^{2}\] \[\tan \theta = \frac{ y }{ x }\] These arethemain ones we need to be aware of. Now, polar coordinates are like: \[(r, \theta)\] So using these conversions, we'll turn our rectangular coordinates into polar coordinates.

OpenStudy (psymon):

So given our point (2, -2), we can start off by finding r. Given our conversion above to get r, I'll use the pythagorean theorem one and do: \[\sqrt{x ^{2}+y ^{2}} = r\] \[\sqrt{(2)^{2}+(-2)^{2}} = \sqrt{8} = 2\sqrt{2}\] So we've determined that r, the first coordinate for our polar coordinate point is 2sqrt2. We good with that? :P

OpenStudy (anonymous):

Si

OpenStudy (psymon):

いいよ! Okay, on to theta!

OpenStudy (psymon):

So the best way is usually to use the conversion: \[\tan \theta = \frac{ y }{ x }\]

OpenStudy (psymon):

So doing that we can say: \[\tan \theta = \frac{ -2 }{ 2 }\] \[\tan \theta = -1\] Now there are two points where tan = -1. So to determine which one we want, we just look at our rectangular coordinates. an x value of 2 and a y value of -2 tells us we are in the 4th quadrant. So in the 4th quadrant, I can say tan = -1 at the angle 7pi/4, which is also 315 degrees. Now apparently, that gives us the answer right away, but let's look at how we would get the 2nd point if you don't mind :3

OpenStudy (anonymous):

So I don't use theta -1 for both?

OpenStudy (psymon):

Well, we'll get to that :P Do you understand how I got the first point, though?

OpenStudy (anonymous):

yeshh

OpenStudy (psymon):

Alright, awesome. So now what the point: \[(2\sqrt{2}, 315)\] means that to get to my point, I need to face 315 degrees and then walk forward 2sqrt2 units. Now there are two ways to get to a point in polar coordinates. We can face the point and walk forward OR we can face 180 degrees in the opposite direction and walk backwards. So if we face the opposite direction, 180 degrees in the other direction, I am then facing 315 - 180 = 135 degrees. SO facing 135 degrees, I now need to walk backwards. Now the radius doesn't change, that's the same, but when you walk backwards, your radius becomes negative. So that being said, your second point is \[(-2\sqrt{2}, 135)\] Kinda make sense? :P

OpenStudy (anonymous):

sorry bout that, yes it does!

OpenStudy (psymon):

Alright, awesome xD

OpenStudy (anonymous):

Find all polar coordinates of point P = (9, 75°). {9,75+2pin}, {-9,75(2n+1)pi} Is this correct?

OpenStudy (psymon):

Seems fine to me, lol.

OpenStudy (anonymous):

okay sweet! what about this one:Find the rectangular coordinates of the point with the polar coordinates (8, 3 divided by 2 pi). (0, -8)

OpenStudy (anonymous):

sorry, I'm trying to get an A in my class and i have a 89 -_-

OpenStudy (psymon):

\[\frac{ 3 }{ 2\pi }?\]

OpenStudy (anonymous):

3/2 (pi)

OpenStudy (psymon):

Lol, that makes more sense then xD

OpenStudy (psymon):

Well, now we just use these conversions: \[x = rcos \theta \] \[y = \sin \theta \] So if theta is 3pi/2, we just need cos and sin at those points and we multiply by 8. SO cos at 3pi/2 is 0, meaning x = 0. sin at 3pi/2 = -1, times 8 means y = -8. So our point is (0, -8) Which itlooks like you got, lol.

OpenStudy (psymon):

Missed the r on the sin conversion, lol. Oh well :P

OpenStudy (anonymous):

hallelujah, THANK YOU

OpenStudy (psymon):

Mhm. These arent bad once you understand whats going on :P

OpenStudy (anonymous):

Yeah, I have my final soon and this is helping a lot since I tend to get distracted easily lol

OpenStudy (psymon):

Lol, I have to multitask. Im playing chess and playing a Japanese game while I do this xD So I get distracted, but I don't get TOO distracted.

OpenStudy (psymon):

If you have your final soon, have you made it to conic sections yet?

OpenStudy (anonymous):

Yes, I wasn't have too much trouble with that. It's just the sequences and stuff

OpenStudy (psymon):

Sequences? You have sequences in your trig work? xD

OpenStudy (anonymous):

It's precalc ): I purposely failed in school so I can retake it online and get an A lol

OpenStudy (psymon):

Odd idea to do, lol. I never had sequences like a lot of people have. So the basic sequences I'm not actually up to speed on. When I did sequences it was in calc 2, so we didn't even do the basic ones, lol.

OpenStudy (psymon):

Well see if you got some other questions ya need. I'll be back in 10-15mins.

OpenStudy (anonymous):

okay thanks

OpenStudy (psymon):

mkay o.o

OpenStudy (anonymous):

Okay found one: lim x^2-2x/(x^4) as x -->0

OpenStudy (psymon):

Interesting. Limits, eh O.o Well, let's see.

OpenStudy (anonymous):

haha do you know them?

OpenStudy (psymon):

Infinity, lol. You'd either say infinity or does not exist, depending on what level we're dealing with. And yeah, I know of them. First topic you see in calculus 1.

OpenStudy (anonymous):

how'd you get that?

OpenStudy (psymon):

Well, again, I don't know if you're expected to put does not exist or infinity, but I'll show you both. And I'm sorry, negative infinity, not infinity xD

OpenStudy (anonymous):

they use DNE

OpenStudy (psymon):

Here is the visual explanation. This is the graph of the function you gave me:  And DNE stands for does not exist.